if x2+px+q=0 and x2+p’x+q’=0 have one root in common, prove that its value is (q-q’)/(p’-p) or (pq’-p’q)/(q-q’)

Grade:11

2 Answers

Vikas TU
6 years ago
Dear Student,
Let common root be m
subtracting the two equations we ge that x = (q - q’)/(p’ - p)
Multiply eq 1 with q' and eq 2 with q
Subtract 2 from 1 to get
M= (p'q-pq') / (q'- q).
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
Charu
5 years ago
If two equation x^2+px+q=0 and other one is x^+p'x+q'=0 then the common root is 
b1c2-b2c1/c1a2 -c2a1 
=pq'-p'/q-q'
 
 
Best of luck 
Charuanurag sharma